Suppose a uniform electric field exists in the room in which you are working, such that the lines of force are horizontal and at right angles to one wall. As you walk toward the wall from which the lines of force emerge into the room, are you walking toward


1
Expert's answer
2021-07-05T16:13:59-0400

Patential higher

"V=\frac{kQ}{r},r<r_1"

Lower potential

"r=r'"

"V=\frac{kQ}{r'}"

Mid potential

r=r1

"V=\frac{kQ}{r_1}"

r increased then potential decrease and

rdecrease and potential increases.